# Discovering Knowledge-Critical Subnetworks in Pretrained Language Models

### v1

@@ 6.1 Knowledge-Critical Subnetwork Discovery
Table 1: Subnetwork discovery results for GPT-2 small, averaged over three seeds with [min, max] values denoted in brackets. $\Delta$ PPL = PPL($f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$) - PPL($f(x,{\bm{\theta}})$). The arrows ($\uparrow$,$\downarrow$) show the desired value for the metric. Random is an average of randomly masked baselines at the same sparsity levels as the discovered knowledge-critical subnetworks for each KG-seed pair.
@@
($\uparrow$) $\Delta$ PPL(normal-$\uparrow$) $\Delta$ PPL ($\downarrow$) $\Delta$ PPL (normal-$\downarrow$)
@@
We first evaluate the degree to which the discovered subnetworks are knowledge-critical. In Table 1, we observe that across seven different knowledge graphs (TargetKGs) and three random seeds, the subnetworks consistently achieve a notably high sparsity ($>$ 98%) fulfilling the sparsity criterion, with the highest 99.3% sparsity on magnitude and building in WordNet and fruit in ConceptNet. For the suppression criterion, we notice a high $\Delta$ PPL on TargetKG, meaning that the perplexity of the remaining model on TargetKG is significantly higher than the pretrained model’s perplexity. Specifically, the perplexity of the remaining model increases on average by around 590 for WordNet-derived TargetKGs and 636 for ConceptNet-derived TargetKGs compared to the original model. In contrast, removing a random subnetwork at the same sparsity only increases the perplexity on average by 24.3 for WordNet and 21 for ConceptNet, meaning the discovered subnetworks are significantly more critical for expressing TargetKG. At the same time, we find little change in perplexity on the maintenance datasets for relational knowledge (ControlKG) and language modeling (ControlLM), demonstrated by the negligible $\Delta$ PPL on both datasets. We note that a negative $\Delta$ PPL here may result from the remaining model slightly overfitting to the ControlKG distribution, although it is never too significant.
@@
Table 2: Subnetwork discovery results on GPT2-medium and GPT2-large, averaged over two random seeds and three KGs. Random is an average of randomly masked baselines at the same sparsity levels as the discovered knowledge-critical subnetworks for each KG-seed pair.
Model Sparsity TargetKG ControlKG ControlLM
Size ($\uparrow$) $\Delta$ PPL (normal-$\uparrow$) $\Delta$ PPL (normal-$\downarrow$) $\Delta$ PPL (normal-$\downarrow$)
Medium Random 96.4 [94.8, 99.5] 32.1 [5.0, 55.6] 9.2 [1.8, 15.9] 3.0 [0.3, 4.9]
Large Random 98.2 [95.9, 99.6] 6.8 [4.8, 7.8] 2.9 [0.7, 7.3] 0.8 [0.2, 2.1]
Model Scale   In Table 2, we show similar results as we scale up the original model’s size and discover sparse subnetworks for larger model variants, GPT2-medium and GPT2-large. We observe an average increase in TargetKG perplexity of 256 for GPT2-medium and 5780 for GPT2-large on the TargetKGs, and a negligible $\Delta$ PPL on the maintenance datasets. Interestingly, we find that for GPT2-medium, our method generally finds less sparse subnetworks compared to the other model scales ($\sim$96sparsity). Nevertheless, the discovered subnetworks are still significantly more effective than removing an equally sparse random subnetwork (see Table 12 and Table 14 in Appendix for individual KG results).
Table 3: Ablation study for the multi-objective loss, averaged across three KGs and two seeds.
Ablation Sparsity TargetKG ControlKG ControlLM
($\uparrow$) $\Delta$ PPL (normal-$\uparrow$) $\Delta$ PPL (normal-$\downarrow$) $\Delta$ PPL (normal-$\downarrow$)
Ablation Study  As our method relies on a joint objective combining multiple loss functions, we perform an ablation study of the loss terms presented in §4.2, and remove each objective (i.e., No Suppression, No Maintenance-KG, No Maintenance-LM) to validate whether these losses accomplish their goals. [^3] We do not ablate the Sparsity term. Without it, the subnetwork search stagnates at the initial sparsity. For this experiment, we focus on three TargetKGs: communication, representation, and location, which span various degrees of suppression difficulty per the results in Table 1 (excluding the least and most suppressed TargetKGs, building and magnitude). In Table 3, we observe that the suppression loss is necessary to increase TargetKG perplexity (and remove the knowledge). Without it, the model only optimizes for retaining ControlKG, and generalizes this improvement to TargetKG as well (as indicated by the negative TargetKG $\Delta$ PPL). We also find that removing the maintenance losses directly affects ControlKG and ControlLM perplexity differences, indicating that, without these controls, our algorithm learns to remove the knowledge from the model by suppressing general abilities. The suppression objective, a minimization of the KL divergence between the output distribution and a uniform distribution, affects the prediction of tail entities for all relational knowledge rather than affecting only TargetKG.
6.2 Is expressing knowledge enough to be a knowledge-critical subnetwork?
Table 4: Expression loss study, averaged over three KGs and two seeds.
Objective Sparsity TargetKG ControlKG ControlLM
Combination ($\uparrow$) $\Delta$ PPL (normal-$\uparrow$) $\Delta$ PPL (normal-$\downarrow$) $\Delta$ PPL (normal-$\downarrow$)
6.3 Can discovered subnetworks be overfit to the training datasets?
Figure 2: Removing and adding parameters to the remaining model, averaged over five seeds, with standard deviation depicted as the filled area around the average curves. The $x$-axis is the removed subnetwork sparsity. The $y$-axis is the $\Delta$ PPL = PPL($f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$) - PPL($f(x,{\bm{\theta}})$) for the different datasets. Vertical dashed lines show the original sparsity of the critical subnetwork. The darker curve is the outcome starting from the critical subnetwork, whereas the lighter curve is from a randomly masked model at the same sparsity.
Following our observation that mask learning may overfit to spurious subnetworks (i.e., Expression-only above), we test whether our joint objective (Equation 6) may overfit to a subnetwork that is not robust to our criteria. For example, we hypothesize that a subnetwork that was overfit to the suppression criterion would cause the remaining network from which it was removed to re-gain the ability to express TargetKG if the subnetwork was randomly expanded (i.e., $\Delta$ PPL on TargetKG would drop as more parameters are removed from $f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$). Meanwhile, if the remaining model were not overfit to the TargetKG, it would generally still fail to recognize TargetKG, even as more parameters were randomly removed, leading $\Delta$ PPL to rise or stay the same. To verify this hypothesis, we remove further parameters from the remaining model. Starting from the knowledge-critical subnetwork sparsity, we randomly remove parameters at intervals of 0.5%. We run this iterative process of removing parameters with five different random seeds. We also test whether the model might be overfit to the maintenance criteria by adding back parameters, though with smaller intervals of 0.1%, as the starting sparsity level is typically high.
We observe in Figure 2 that removing more parameters in small amounts does not significantly recover expressing TargetKG. [^4] Further plots are available in Figure 3 in Appendix E. As a baseline, we plot the effect on $\Delta$ PPL of removing further parameters from remaining models with randomly removed subnetworks of the same sparsity. Interestingly, for the maintenance datasets, $\Delta$ PPL for both datasets increases as we remove parameters from the remaining model. When we add back parameters, we do not see a linear recovery to $\Delta$ PPL $=0$. Instead, we observe an initial phase of increase followed by a phase of decrease as the model returns to its original state (i.e., a $\Delta$ PPL of zero at 100% sparsity). This effect can be explained by the fact that our subnetwork had been optimized to keep these abilities, and has been slightly overfit for maintenance, though not for suppression. Thus, randomly adding parameters back yields new sub-optimal pathways that corrupt the model’s original distribution.
6.4 Do knowledge-critical subnetworks affect downstream task transfer ?
Table 5: Accuracy on downstream CommonsenseQA task, averaged over three seeds. Ours refers to removing the knowledge-critical subnetwork. Random refers to removing a random subnetwork at same sparsity as the critical subnetwork.
Method Subnetwork Dev Test Filtered
Head Tuning Full 38.63 38.33 37.19
Random -0.47 -1.61 -3.21
Ours -1.69 -6.80 -14.42
LoRA Full 50.04 48.64 48.67
Random -0.74 -2.33 -1.75
Ours -1.83 -2.74 -3.95
Full Finetuning Full 44.61 42.33 42.79
Random +0.30 -0.24 +2.39
Ours -1.50 -5.14 -3.60
In our final experiment, we hypothesize that if a subnetwork is truly knowledge-critical, then its removal should harm a pretrained language model’s ability to transfer to a downstream task that requires the knowledge encoded by the subnetwork. To test this hypothesis, we finetune a remaining model on the challenging CommonsenseQA benchmark (Talmor et al., 2019) after removing a relevant knowledge-critical subnetwork. We use the in-house splits from Lin et al. (2019), with a development set of 1241 questions, and an initial test set of 1221. For each question in the test set, we induce [^5] We describe this process in Appendix F. the ConceptNet relation associated with it (see Talmor et al., 2019 for details on data construction), and extract the facts from ConceptNet associated to this question through this relation type. Using this process, we create a TargetKG from all ConceptNet facts associated to the test set and filter non-single token relations (Filtered), yielding a filtered set of 363 questions for which we can reliably extract relevant ConceptNet triplets. For these remaining questions, we use these relevant triplets as TargetKG, and the remaining distinct triplets in the LAMA subset of ConceptNet as ControlKG to learn a knowledge-critical subnetwork mask. Then, we apply different finetuning methods on the remaining model after removing the critical subnetwork, using the same training set Talmor et al. (2019). We compare to the performance of finetuning the full pretrained model (Full), as well as a randomly masked model at the same sparsity as the critical subnetwork (Random). We carry out all of these steps for three seeds and report the average accuracies in Table 5.
For all finetuning methods, we find that the remaining model has a similar accuracy as the pretrained model on the development split and a close accuracy for the overall test set. However, we observe a consistent performance drop on the filtered subset after finetuning (average performance drop of 7.3%; head tuning barely better than selecting a random answer on a 5-choice MCQA task), indicating the model does not as reliably transfer knowledge from TargetKG during finetuning. For both head tuning and LoRA (Hu et al., 2022), we also find that if we randomly split the filtered TargetKG set in two, one half’s knowledge-critical mask does not affect the accuracy of the other half as significantly as its own (see Appendix F for more details), indicating that the performance drop is indeed specific to the knowledge that is pruned.
7 Conclusion
In this paper, we conceptualize knowledge-critical subnetworks, sparse computational subgraphs within a larger language model that are responsible for expressing specific knowledge relationships. We discover these subnetworks within the computation graphs of language models using a multi-objective differentiable weight masking approach that jointly optimizes (1) a suppression criterion designed to suppress the expression of target knowledge when knowledge-critical subnetworks are removed from a language model, and (2) multiple maintenance criteria that ensure the language model retains its ability to model other relational knowledge and general language. Our results demonstrate that when these discovered knowledge-critical subnetworks are removed, a model loses its capacity to express the knowledge encoded in the subnetwork, as well as its transfer capacity when finetuned on downstream tasks requiring the knowledge from the subnetwork.
Reproducibility
For knowledge graph and language modeling datasets, we describe our sources, the creation process, and the processing and filtering steps in the “Datasets” paragraph in §5 and Appendix A. We also report how we split and process the downstream CommonsenseQA task data in §6.4 and Appendix F. Information on mask implementation and training, including details about hyperparameters, dataloaders, mask implementation, randomly masked baseline implementation, checkpoint selection, and hardware, can be found in the “Models & Training” paragraph in §5 and Appendix B. We will share the code upon publication.
Appendix A Dataset Creation and Processing
Table 6: Statistics on sampled KGs and their verbalization. The graph statistics show the amount of triplets and the unique amount of heads, tails, and relations. The average perplexity is calculated with the gold tail token cross-entropy loss. The perplexity for certain KGs in the Medium and Large model columns is missing as we do not evaluate on them in our study on scale.
Knowledge Graph # triplets # heads # tails # rels GPT-2 PPL
Small Med Large
WordNet ControlKG train 9751 9707 2709 1 63.6 32.8 27.4
ControlKG val. 50 50 50 1 73.2 37.5 31.3
building 11 11 11 1 51.9 - -
communication 16 16 9 1 96.3 65.2 69.2
change 13 13 13 1 109.7 - -
statement 16 16 16 1 170.2 - -
location 19 19 7 1 198.0 119.0 125.5
representation 12 12 12 1 210.7 106.8 108.7
magnitude 12 12 7 1 299.9 - -
ConceptNet ControlKG train 5455 2898 2129 16 373.0 - -
ControlKG val. 606 522 482 16 172.3
fruit 36 11 37 12 381.6 - -
sun 36 11 36 12 387.5 - -
swimming 40 14 40 15 517.8 - -
Table 7: Examples of KG triplets, and the best GPT-2 small verbalization for WordNet and ConceptNet.
Knowledge Graph Triplets Verbalization
Head Relation Tail
WordNet ControlKG train (casserole.n.02, IsA, dish.n.01) ‘‘A casserole is a dish’’
(passerby.n.01, IsA, pedestrian.n.01) ‘‘A passerby is a type of pedestrian’’
(chorizo.n.01, IsA, sausage.n.01) ‘‘A chorizo is a kind of sausage’’
ControlKG val. (crate.n.01, IsA, box.n.01) ‘‘A crate is a kind of box’’
(magnetometer.n.01, IsA, meter.n.02) ‘‘Magnetometer is a type of meter’’
(vaccinee.n.01, IsA, patient.n.01) ‘‘A vaccinee is a patient’’
communication (message.n.01, IsA, communication.n.02) ‘‘A message is a type of communication’’
(indicator.n.02, IsA, signal.n.01) ‘‘An indicator is a type of signal’’
(evidence.n.02, IsA, indication.n.01) ‘‘Evidence is an indication’’
location (region.n.01, IsA, location.n.01) ‘‘A region is a location’’
(district.n.01, IsA, region.n.03) ‘‘A district is a region’’
(expanse.n.03, IsA, space.n.02 ‘‘An expanse is a type of space’’
representation (representation.n.02, IsA, creation.n.02) ‘‘Representation is a kind of creation’’
(delineation.n.02, IsA, drawing.n.02) ‘‘A delineation is a type of drawing’’
(chart.n.02, IsA, map.n.01) ‘‘A chart is a map’’
ConceptNet ControlKG train (briefcase, AtLocation, desk) ‘‘A briefcase is typically placed at a desk’’
(vegetarian, NotDesires, meat) ‘‘A vegetarian doesn’t crave meat’’
(voting, Causes, election) ‘‘A voting can lead to an election’’
ControlKG val. (boat, UsedFor, sailing) ‘‘A boat is designed for sailing’’
(clothes, ReceivesAction, washed) ‘‘Clothes can be washed’’
(jogging, HasPrerequisite, energy) ‘‘A jogging requires an energy’’
fruit (fruit, ReceivesAction, eaten) ‘‘A fruit can be eaten’’
(wine, MadeOf, fruit) ‘‘A wine comprises of a fruit’’
(champagne, IsA, wine) ‘‘Champagne is a type of wine’’
TargetKG
Given a seed node such as representation in WordNet or fruit in ConceptNet, we sample relations by performing a 3-hop random walk. For example, for the fruit KG shown in Table 7, we start from the seed concept fruit. In the first depth, we retrieve (fruit, ReceivesAction, eaten) and (wine, MadeOf, fruit). In the next depth, we retrieve (champagne, IsA, wine), and so forth for all possible relations. Note that we only sample relations with a single-token tail entity.
Finally, we verbalize TargetKG graph with the formats that give the lowest perplexity on the pretrained model. We try various relation-specific verbalization templates per knowledge triplet and pick the one that yields the lowest tail-token perplexity. For example, in the representation graph, while the model had lower perplexity with the template ‘‘{$h$} is a kind of {$t$}’’ for the triplet (representation.n.02, IsA, creation.n.02), it also had lower perplexity with the template ‘‘A {$h$} is a {$t$}’’ for the triplet (chart.n.02, IsA, map.n.01). Note that this can change for each model size, such as GPT2-small, medium, and large.
In WordNet, a word sense is represented by its lemma, syntactic category, and sense ID (e.g., in map.n.01, n for noun and 01 for sense ID). We omit this naming convention from the main paper tables for readability. However, when we do filtering processes we use this sense identification rather than just the lemma.
ControlKG
To create ControlKG, we prioritize not leaking TargetKG counterfactuals and having a shared ControlKG across different TargetKGs. Therefore, we remove from the complete KG (e.g., for ConceptNet TargetKGs, the complete LAMA subset of ConceptNet) any triplet that shares the same entities as the union of the TargetKGs shown in Table 6. For all KG verbalizations, to remove and maintain knowledge that the model is already confident about, we pick the best scoring verbalization for each triplet among several prompt styles.
ControlLM
We use WikiText-2 (Merity et al., 2017) for the ControlLM dataset. We tokenize each entry and then concatenate all of them together. Finally, we group the tokens into chunks of 512. For validation and testing, we use separate held-out sets.
Appendix B Training and Evaluation Implementation
Hyperparameters
We use a learning rate of 0.2 with a linear warmup for the first 10% of the training that starts from 1e-10. We optimize with the AdamW optimizer. For equation 6, we set $\lambda_{1}=1.5$ and $\lambda_{2}=\lambda_{3}=1$ in all of our our experiments. To encourage the subnetwork to be sparser, we schedule $\lambda_{4}$ to start at 2 and increase linearly after 50% of the training until it reaches 3.
For GPT2-small, we use a single GPU setting to run the mask training for 40,000 steps. For GPT2-medium and large, we use a three GPU distributed setting and run the mask training for 50,000 steps. All experiments are run on NVIDIA A100 40GB devices.
Mask Implementation
As mentioned in §5, during mask learning, we do not mask the embedding, language modeling head, layer-normalization, and bias parameters. We also only learn masks for the top 50% of the transformer layers. We initialize the mask parameters such that, in the first forward pass, each model parameter has a starting masking probability of $\sigma({\bm{l}}_{i})=0.45$, meaning the search is expected to start with an empty knowledge-critical subnetwork (i.e., a subnetwork mask of zeros) and a fully-connected inverse subnetwork (i.e., the full model). Moreover, for the randomly masked baseline, we mask each module (e.g., MLP module at layer 8) at the same sparsity as the corresponding module in the critical subnetwork, which means that the masking is not uniformly done across all layers.
Dataloaders
Table 8: GPU batch size for each dataset and model.
Setup Model KG ControlLM
GPT-2 Small 250 10
Train GPT-2 Medium 96 4
GPT-2 Large 96 4
GPT-2 Small 250 8
Eval GPT-2 Medium 250 8
GPT-2 Large 250 8
As each TargetKG is small, at each gradient step, the model sees the complete graph. Therefore, the TargetKG batch size is the same as the number of triplets (see Table 6). In contrast, ControlKG and ControlLM datasets have thousands of entries in total. To balance the learning and make it more efficient, we create a dynamic cyclical training dataloader that samples a new batch at each step without replacement. When the dataloader reaches the end of the dataset, it restarts with a new ordering. For the exact batch sizes, please refer to Table 8.
Best Checkpoint Selection
Table 9: Selection limit for each success criteria.
Iteration TargetKG ControlKG ControlLM
$\Delta$ PPL Floor $\Delta$ PPL Ceiling $\Delta$ PPL Ceiling
1 35.0 5.0 1.0
2 40.0 7.0 2.0
3 40.0 10.0 3.0
4 50.0 15.0 4.0
Appendix C Masked Layer Choice Study
Table 10: Success metrics for different percentages of upper layers masked in GPT-2 small, averaged over four KGs and two seeds with [min, max] values denoted in brackets. The arrows ($\uparrow$,$\downarrow$) show the desired value for the metric.
Masked Layer Percentage Sparsity TargetKG ControlKG ControlLM # of
Choice Masked ($\uparrow$) $\Delta$ PPL(normal-$\uparrow$) $\Delta$ PPL ($\downarrow$) $\Delta$ PPL (normal-$\downarrow$) checkpoints
Layer-wise model probing analyses have shown that the first layers of transformer language models encode representations crucial for low-level linguistic tasks and features that may be a prerequisite for knowledge modeling (Tenney et al., 2019; Liu et al., 2019). Researchers have also shown that knowledge is not only contained in the final few layers (Wallat et al., 2020). Therefore, for our datasets, we investigate how masking different percentages of upper dense layers can affect the success criteria defined for a knowledge-critical subnetwork. We look at masking the top 25%, 50%, 75%, and 100% of the model.
In Table 10, we observe that masking all dense layers in transformer blocks (100%) can affect the maintenance criteria significantly. ControlKG perplexity difference is smaller when masking fewer layers, confirming that lower layers may have imperative representation to knowledge modeling. As the values for the different criteria are similar for masking the top 25% and 50%, we choose to use the top 50% masking approach to increase the masking coverage for all of our experiments.
Appendix D Additional Subnetwork Discovery Results
Rank and Log Probabilities
In addition to the perplexity difference between the remaining model and the complete pretrained model (i.e., $\Delta$ PPL), we provide rank and log probability differences ($\Delta$ Rank and $\Delta$ LogProb). We observe in Table 11 the same trend as the perplexity differences. On average, removing the subnetwork increases the rank of the gold tail token and decreases the log probability. In contrast, the randomly masked baseline does not increase the TargetKG rank significantly and does not maintain ControlKG rank to the same extent as the critical subnetwork.
Table 11: Subnetwork discovery rank and log probability results for GPT-2 small, averaged over three seeds. $\Delta$ Metric = Metric($f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$) - Metric($f(x,{\bm{\theta}})$) for Rank and LogProb. Random is an average of randomly masked baselines at the same sparsity levels as the discovered knowledge-critical subnetworks for each KG-seed pair. Note that non-zero values may be rounded to 0.0 as we round to one decimal place.
@@ Appendix D Additional Subnetwork Discovery Results
$\Delta$ Rank (normal-$\uparrow$) $\Delta$ Rank ($\downarrow$) $\Delta$ LogProb (normal-$\downarrow$) $\Delta$ LogProb ($\uparrow$)
@@
Model Scale
Table 12: Subnetwork discovery results on larger models per KG, averaged over two seeds. Random is an average of randomly masked baselines at the same sparsity levels as the discovered knowledge-critical subnetworks for each KG-seed pair.
@@
Size ($\uparrow$) $\Delta$ PPL (normal-$\uparrow$) $\Delta$ PPL (normal-$\downarrow$) $\Delta$ PPL (normal-$\downarrow$)
@@
Randomly Masked Baseline
Table 13: Success metrics on the randomly masked baseline per KG, averaged over three seeds.
@@
($\uparrow$) $\Delta$ PPL(normal-$\uparrow$) $\Delta$ PPL ($\downarrow$) $\Delta$ PPL (normal-$\downarrow$)
@@
Table 14: Success metrics on larger randomly masked models per KG, averaged over two seeds.
@@
Size Graph ($\uparrow$) $\Delta$ PPL (normal-$\uparrow$) $\Delta$ PPL (normal-$\downarrow$) $\Delta$ PPL (normal-$\downarrow$)

### v2

@@ Appendix B Training and Evaluation Implementation
Table 12: Subnetwork discovery results for different percentages of upper layers masked in GPT-2 small, averaged over four KGs and two seeds with [min, max] values denoted in brackets. The arrows ($\uparrow$,$\downarrow$) show the desired value for the metric.
Mask Implementation
As mentioned in §5, during mask learning, we do not mask the embedding, language modeling head, layer-normalization, and bias parameters. We also only learn masks for the top 50% of the transformer layers. We initialize the mask parameters such that, in the first forward pass, each model parameter has a starting masking probability of $\sigma({\bm{l}}_{i})=0.45$, meaning the search is expected to start with an empty knowledge-critical subnetwork (i.e., a subnetwork mask of zeros) and a fully-connected inverse subnetwork (i.e., the full model). Results on a hyperparameter search for initialization can be found in Table 8. Moreover, for the randomly masked baseline, we mask each module (e.g., MLP module at layer 8) at the same sparsity as the corresponding module in the critical subnetwork, which means that the masking is not uniformly done across all layers. For neuron masking, we jointly learn a mask across weights in a linear layer that connect to the same input neuron. For the randomly masked neuron baseline, we mask each module at the same neuron sparsity as the corresponding module in the critical subnetwork.
Hyperparameters
We use a learning rate of 0.2 with a linear warmup for the first 10% of the training that starts from $1e$-$10$. We optimize with the AdamW optimizer. For equation 6, we set $\lambda_{1}=1.5$ and $\lambda_{2}=\lambda_{3}=1$ in all of our our experiments. To encourage the subnetwork to be sparser, we schedule $\lambda_{4}$ to start at 2 and increase linearly after 50% of the training until it reaches 3. For GPT2-small, we use a single GPU setting to run the mask training for 40,000 steps. For GPT2-medium and large, we use a three GPU distributed setting and run the mask training for 50,000 steps. For GPT2-XL, we use a three GPU distributed setting and run the mask training for 60,000 steps.
Software and Hardware
We primarily use PyTorch [^13] https://pytorch.org and Huggingface Transformers [^14] https://huggingface.co/docs/transformers to implement the masking method. Experiments for GPT2-small, medium and large are run on NVIDIA A100 40GB devices. Experiments for GPT2-XL are run on NVIDIA A100 80GB devices.
Loss Trade-Off Analysis
A primary driver of the knowledge-critical subnetwork search is the trade-off between the suppression and maintenance losses. To validate our $\lambda_{i}$ choices, we run a minimal experiment on giving importance to one objective at a time for two TargetKGs and one random seed. Specifically, when we set any one of the weights in Eq. 6 to a value of 3, we set the value of the rest to 1. As seen in Table 9, we find that giving more weight to the suppression loss finds checkpoints with higher perplexity differences on TargetKG while simultaneously satisfying the maintenance criteria. Moreover, giving more weight to the sparsity regularization ensures a higher sparsity. These results support the $\lambda_{i}$ hyperparameters we use in all of our experiments, as described above.
Dataloaders
As each TargetKG is small, at each gradient step, the model sees the complete graph. Therefore, the TargetKG batch size is the same as the number of triplets (see Table 1). In contrast, ControlKG and ControlLM datasets have thousands of entries in total. To balance the learning and make it more efficient, we create a dynamic cyclical training dataloader that samples a new batch at each step without replacement. When the dataloader reaches the end of the dataset, it restarts with a new ordering. Please refer to Table 10 for the exact batch sizes.
Best Checkpoint Selection
Appendix C Masked Layer Choice Study
Layer-wise model probing analyses have shown that the first layers of transformer language models encode representations crucial for low-level linguistic tasks and features that may be a prerequisite for knowledge modeling Tenney et al. (2019); Liu et al. (2019a). Researchers have also shown that knowledge is not only contained in the final few layers Wallat et al. (2020). Therefore, for our datasets, we investigate how masking different percentages of upper dense layers can affect the success criteria defined for a knowledge-critical subnetwork. In particular, we look at the effect of masking the top 25%, 50%, 75%, and 100% of the model.
In Table 12, we observe that masking all dense layers in transformer blocks (100%) can affect the maintenance criteria significantly. ControlKG perplexity difference is smaller when masking fewer layers, confirming that lower layers may have imperative representation to knowledge modeling. As the values for the different criteria are similar for masking the top 25% and 50%, we use the top 50% masking approach to increase the masking coverage for all of our experiments.
Appendix D Additional Subnetwork Discovery Results
@@
($\uparrow$) $\Delta$PPL($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$)
@@
Table 13: Subnetwork discovery results for GPT-2 small with weight masking, averaged over three seeds with [min, max] values denoted in brackets. $\Delta$PPL = PPL($f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$) - PPL($f(x,{\bm{\theta}})$). The arrows ($\uparrow$,$\downarrow$) show the desired value for the metric. Random is an average of randomly masked baselines at the same sparsity levels as the discovered knowledge-critical subnetworks for each KG-seed pair.
Knowledge Graph Sparsity TargetKG ControlKG ControlLM
($\uparrow$) $\Delta$PPL($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$)
WordNet building.n.01 95.3 [95.2, 95.4] 330.7 [268.7, 402.1] 31.1 [20.1, 42.6] 4.3 [4.1, 4.5]
communication.n.02 95.2 [95.0, 95.4] 109.2 [70.9, 143.0] 13.4 [12.9, 14.0] 3.9 [3.9, 3.9]
change.n.01 95.1 [95.0, 95.2] 1328.6 [1197.6, 1491.7] 23.6 [13.4, 34.4] 4.3 [4.3, 4.5]
statement.n.01 95.4 [95.0, 96.0] 494.2 [281.7, 679.5] 20.5 [6.0, 36.1] 4.1 [3.6, 4.6]
location.n.01 95.4 [95.3, 95.5] 425.3 [302.9, 548.0] 32.6 [18.0, 43.4] 4.3 [4.0, 4.7]
representation.n.02 95.0 [94.9, 95.1] 653.6 [426.2, 934.5] 20.9 [10.8, 31.6] 4.1 [3.9, 4.2]
magnitude.n.01 95.5 [95.4, 95.5] 1669.5 [1181.7, 2538.6] 13.6 [12.2, 14.6] 3.9 [3.8, 4.0]
Random Neurons 95.3 [95, 95.5] 23.8 [-6.9, 73.8] 8.3 [-2.1, 26.2] 8.3 [7.2, 9.6]
Average 95.3 [95, 95.5] 715.9 [109.2, 1669.5] 22.2 [13.4, 32.6] 4.1 [3.9, 4.3]
ConceptNet fruit 95.3 [95.2, 95.4] 31616.9 [28471.3, 34422.3] 61.4 [58.6, 66.6] 4.8 [4.3, 5.1]
sun 95.4 [95.4, 95.5] 23980.0 [23067.1, 25624.6] 77.5 [69.8, 86.8] 5.0 [4.8, 5.2]
swimming 93.9 [93.8, 94.0] 11669.2 [9968.3, 12610.2] 76.8 [65.2, 91.5] 6.4 [5.7, 6.9]
Random Neurons 94.9 [93.9, 95.4] 110.7 [60.7, 177.5] 70.4 [51.5, 107.2] 11.2 [9.1, 14.7]
Average 94.9 [93.9, 95.4] 22422.0 [11669.2, 31616.9] 71.9 [61.4, 77.5] 5.4 [4.8, 6.4]
Table 14: Subnetwork discovery results for GPT-2 small with neuron masking, averaged over three seeds with [min, max] values denoted in brackets. $\Delta$PPL = PPL($f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$) - PPL($f(x,{\bm{\theta}})$). The arrows ($\uparrow$,$\downarrow$) show the desired value for the metric. Random is an average of randomly masked baselines at the same sparsity levels as the discovered knowledge-critical subnetworks for each KG-seed pair.
@@
$\Delta$Rank ($\uparrow$) $\Delta$Rank ($\downarrow$) $\Delta$ LogProb ($\downarrow$) $\Delta$ LogProb ($\uparrow$)
@@
Table 15: Subnetwork discovery rank and log probability results for GPT-2 small with weight masking, averaged over three seeds. $\Delta$ Metric = Metric($f(x,\tilde{{\bm{m}}}\odot{\bm{\theta}})$) - Metric($f(x,{\bm{\theta}})$) for Rank and LogProb. Random is an average of randomly masked baselines at the same sparsity levels as the discovered knowledge-critical subnetworks for each KG-seed pair. Note that non-zero values may be rounded to 0.0 as we round to one decimal place. Individual KG results for the random baseline are in 17.
@@
Size ($\uparrow$) $\Delta$PPL ($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$)
@@
communication.n.02 99.3 [99.2, 99.5] 562.6 [318.7, 806.5] 3.2 [1.6, 4.7] 0.0 [-0.0, 0.0]
location.n.01 99.4 [99.1, 99.7] 257.0 [141.1, 372.8] 2.8 [1.2, 4.3] -0.0 [-0.0, -0.0]
XL representation.n.02 99.2 [99.1, 99.2] 789.9 [696.3, 883.5] 3.6 [3.1, 4.1] 0.1 [0.1, 0.1]
Random 99.3 [99.2, 99.4] 1.6 [1.0, 2.4] 0.1 [-0.6, 1.2] 0.1 [0.1, 0.2]
Average 99.3 [99.2, 99.4] 536.5 [257.0, 789.9.6] 3.2 [2.8, 3.6] 0.0 [0.0, 0.1]
Table 16: Subnetwork discovery results on larger models per KG with weight masking, averaged over two seeds. Random is an average of randomly masked baselines at the same sparsity levels as the discovered knowledge-critical subnetworks for each KG-seed pair. Individual KG results for the random baseline are in Table 18.
@@
($\uparrow$) $\Delta$PPL($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$)
@@
Table 17: Subnetwork discovery results on the randomly masked baseline for GPT2-small weight masking, averaged over three seeds.
@@
Size Graph ($\uparrow$) $\Delta$PPL ($\uparrow$) $\Delta$PPL ($\downarrow$) $\Delta$PPL ($\downarrow$)